Please help me!!
directions: Identify the conic section represented by the eqation by writing the equation in standard form.
(-x^2) + (y^2) + (4y) - (16) = 0

thank you

Answer by stanbon(29603)

(-x^2) + (y^2) + (4y) - (16) = 0
y^2+4y+4 - x^2 = 16+4
[(y+2)^2/20 - x^2/20 =1
Hyperbola because of the negative on the x^2.
=========
Cheers,
Stan H.